# PEVLM: Parallel Encoding for Vision-Language Models

### v2

@@ Introduction
In recent years, Vision-Language Models (VLMs) have emerged as a central research focus at the intersection of computer vision and natural language processing, demonstrating impressive performance in multimodal understanding and generation tasks (Alayrac et al. 2022; Li et al. 2020; Radford et al. 2021; Zhu et al. 2023). As the capabilities of these models continue to strengthen, their application scenarios have also expanded, including robotics (Black, Lee et al. 2024; Cheang, Huang et al. 2024; Prasad, Singh et al. 2024), autonomous driving (Gao, Lin et al. 2024; Hu, Wu et al. 2023; Wang, Zhao et al. 2024), and healthcare (Liu, Zhang et al. 2024). Correspondingly, the length of the videos has also increased significantly.
However, applying VLMs to long-video data introduces a significant computational challenge: The quadratic complexity of transformer-based attention mechanisms causes computation to scale rapidly with the input sequence length (Beltagy, Peters, and Cohan 2020; Vaswani et al. 2017), especially during the prefilling stage, where all input tokens must be processed before generating any output. This severely limits the scalability of VLMs in long-video understanding tasks. In practical scenarios where the entire video must be encoded before decoding begins—such as video captioning, retrieval, or reasoning—this inefficiency becomes a critical bottleneck. Therefore, optimizing the prefilling phase is essential to efficiently deploy VLMs on long video tasks.
A widely adopted technique in LLMs to speed up long context prefilling is the parallel encoding mechanism (Li et al. 2024b; Liu et al. 2024; Ma, Wang, and Tian 2025; Ratner et al. 2023; Yang, Chen, and Chen 2025; Yen, Gao, and Chen 2024; Lu et al. 2024). In these approaches, the input context is divided into blocks, and each block is encoded into Key-Value (KV) states independently. This means that tokens from different blocks do not attend to each other. Since the parallel encoding mechanism omits attention computation across blocks, it reduces the computational complexity from $O(L^{2})$ to $O(L)$. The parallel encoding mechanism also mitigates the problem of "lost in the middle" (Liu et al. 2023), and in long-context scenarios, it can even achieve accuracy surpassing that of Full-Attention (Yang, Chen, and Chen 2025).
@@
In addition, there are several widely discussed methods such as Native Sparse Attention (Yuan et al. 2025) from Deepseek and MoBA (Lu et al. 2025), which adopt similar algorithmic principles and achieve acceleration effects comparable to parallel encoding. However, both approaches involve modifications to the model architecture and thus require retraining from scratch. This poses a challenge for many inference platforms, which typically do not have control over users’ model preferences. The issue is particularly pronounced in domains such as autonomous driving, where a reliable models are often continuously trained and iterated over years. Consequently, we focus on finetune-free optimization methods to ensure compatibility and ease of deployment.
Given all of these, a key question naturally arises: Can existing methods be directly applied to VLMs to accelerate inference and deployment pipelines? In Table 1, we evaluate several models with existing parallel encoding methods on LongVideoBench (Zhang and Wang 2024), VideoMME (Fu et al. 2024), EgoSchema (Mangalam, Akshulakov, and Malik 2023) and MVBench (Li et al. 2024a) benchmarks. Although existing parallel encoding methods achieve some effectiveness in shorter videos, we observe varying degrees of accuracy degradation across different models on longer videos, particularly for Qwen2.5-VL. To investigate this issue, we analyze the attention weight distributions under parallel encoding and identify a misalignment between Full-Attention and parallel encoding. Existing methods developed for LLMs fail to effectively address this misalignment in VLMs. Building upon this observation. We conducted a series of studies and experiments and ultimately propose a novel attention mechanism for VLMs, termed Parallel Encoding for Vision-Language Models (PEVLM). PEVLM is a fine-tuning-free method specifically designed to accelerate the prefilling stage in long-video processing with VLMs. Our contributions involve:
@@
  • We propose PEVLM to recover the accuracy of parallel encoding by applying three targeted alignment steps: (i) Segmenting contexts into blocks by video frames rather than tokens. (ii) Using the system promots and the initial frames of the video as Sink Block for all blocks to avoid the duplication of abnormal distribution of initial tokens. (iii) Preserving the sequential position embeddings instead of resuing position embeddings across blocks. PEVLM achieves superior accuracy compared to other parallel-encoding-based methods on long video processing tasks, with up to a 9.73% improvement across different models.
  • PEVLM reduces the computational complexity of the attention mechanism from $O((T\times N)^{2})$ to $O(T\times N)$, which leading to a significant improvement in prefilling efficiency. (i) For 100k token contexts prefilling, the attention layer achieves a 7.47× speedup, while the end-to-end 2.58x speedup without compromising generation quality. This highlights the practical viability and superiority of PEVLM in cloud deployment scenarios. (ii) Under fixed latency constraints, PEVLM increases accuracy from 23.26% to 61.03%, showcasing its critical value for latency sensitive applications.
@@
Figure 1: We collected the distribution of attention weights under different scenarios and computed their moving averages for better observation. The top row shows the attention weights for all tokens, while the bottom row shows those for the video tokens. (a) Full-Attention: The textual tokens and the initial video frames receive the highest attention weights, whereas the remaining video frames are assigned relatively lower attention weights. (b) Parallel Encoding (Reusing Position Embedding): The overall attention weights distribution resembles Full-Attention, with higher attention weights for the textual tokens and initial video frames. While compared to Full-Attention, the initial video frames receive even higher attention weights, and the remaining video frames exhibit more uniform and lower attention weights. (c) Parallel Encoding (Sequential Position Embedding): The overall attention weights distribution remains similar to (a) and (b). However, the attention weight distribution within the video frames more closely aligns with (a)).
@@
(i) As shown in Figure 1(a), initial segments of both text and video receive high attention weights, emphasizing the need to include early video frames when selecting sink tokens.
@@
These observations yield two key insights for applying parallel encoding to VLMs: The shared sink block in parallel encoding should include early video frames in addition to textual prompts, as they function as important attention sinks. Maintaining the sequential position embeddings of video tokens is essential for preserving temporal consistency; reusing position embeddings across blocks can disrupt temporal alignment. These findings underscore the need for a VLM-specific parallel encoding strategy, which we introduce in the next section.
@@
With the observation in the last section, we design PEVLM to improve computational efficiency. PEVLM enables a seamless shift to parallel encoding without requiring training while maintaining most of the model’s capabilities. Our approach adaptively aligns the distribution of attention weights between Full-Attention and parallel encoding, thereby boosting efficiency and performance.
@@
In a standard Softmax attention, we attend the query to all past KV states using the following formula:
@@
Attention Sink. First discovered by StreamingLLM (Lab and AI 2023), that one or more tokens at the start position of an LLM receive higher attention scores. APE (Yang, Chen, and Chen 2025) and Star-Attention (Liu et al. 2024) have demonstrated that sharing a common attention sink across parallel encoding blocks significantly enhances model accuracy. Some VLM-related studies also show that the attention sink phenomenon exists in VLMs but is more distributed among image tokens (Kang et al. 2025).
Position Embedding. To effectively process sequential input, LLMs require position embeddings, such as absolute position embeddings (Devlin et al. 2019; Vaswani et al. 2017) and relative position embeddings (Press, Smith, and Lewis 2022; Su et al. 2024). For VLMs, contextual positional information is more critical. In particular, for some VLM models like Qwen2.5-VL, their position embeddings also incorporate spatio-temporal information between video frames, making position embeddings increasingly essential for VLMs.
@@
Computational Complexity. In standard softmax attention mechanisms, the total number of operations is given by
@@
Sink Size & Block Size
In addition, as analyzed in the last section, the number of frames selected for the Sink Block depends primarily on the distribution of attention weights. The optimal strategy is to include all tokens with significantly higher attention scores at the beginning of the video in the Sink Block. However, as shown in Equation 11, a larger sink size results in higher latency, leading to a trade-off between accuracy and efficiency. A similar trade-off also arises in the selection of block size.
@@
For different blocks, PEVLM applies distinct computation methods to ensure efficient processing and information integration:
$\text{Attn}_{s}\ =f(Q_{s},K_{s},V_{s}),$ (3)
@@
where $s$ and $q$ are the Sink Block and question block, $c_{i}$ is the ith Context Blocks; $Attn_{s}$, $Attn_{c_{i}}$ and $Attn_{q}$ denote the attention of the $s$, $c_{i}$ and $q$. $f(\cdot)$ represents the standard Softmax attention in Equation (1), and $Q$, $K$ and $V$ denote the query, key, and value matrices.
The total OP of PEVLM is
$OP_{\text{PEVLM}}=OP_{\text{Sink}}+N\times OP_{\text{ContextBlock}}+OP_{\text{Quest}},$ (6)
$OP_{\text{Sink}}=2HS^{2},$ (7)
$OP_{\text{ContextBlock}}=2HB(S+B),$ (8)
$OP_{\text{Quest}}=2HQL,$ (9)
where $S$ denotes the sink size, $N$ denotes the block number, $B$ denotes the block size, $Q$ denotes the query size and $H$ is the hidden size. $L$ is the total token number and $L=S+N\times B+Q$.
@@
As the block size is a static number, the computational complexity of PEVLM is $O(L)$.
Model Method MVBench EgoSchema VideoMME LongVideoBench
$17s\ in\ avg.$ 3 mins in avg. $<2mins$ $[2mins,1h]$ $<1min$ $[1min,1h]$
Qwen2.5-VL- Full Attn 68.47% 57.00% 74.56% 56.72% 72.85% 55.33%
7B-Instruct Block Attn 66.03% 47.40% 65.22% 50.39% 70.64% 47.85%
Max pixels: 90M APE (T=1.0) 66.03% 57.20% 50.56% 1.06% 54.02% 3.38%
Tokens: 100k Star Attn 60.19% 19.80% 3.00% 0.00% 49.03% 0.00%
PEVLM 68.44% 61.00% 74.33% 58.83% 74.52% 57.58%
LongVILA- Full Attn 61.92% 57.00% 66.33% 49.39% 65.10% 47.44%
7B-256f Block Attn 57.81% 59.00% 64.89% 51.44% 61.77% 48.05%
Frames: 256 APE (T=1.0) 60.00% 60.60% 68.56% 51.61% 63.71% 46.93%
Tokens: 66k Star Attn 63.39% 61.60% 70.11% 54.17% 63.43% 47.44%
PEVLM 63.53% 61.40% 69.44% 54.33% 65.93% 49.18%
LLaVA-Video- Full Attn 60.67% 58.20% 75.33% 59.99% 72.02% 55.53%
7B-Qwen2 (128) Block Attn 56.86% 51.60% 67.56% 55.50% 63.99% 52.36%
Frames: 128 APE (T=1.0) 58.00% 57.60% 72.22% 58.19% 67.87% 54.61%
Tokens: 26k Star Attn 59.31% 57.80% 72.89% 58.71% 68.98% 54.92%
PEVLM 60.00% 58.20% 74.33% 59.99% 72.02% 55.53%
Position Encoding
Figure 3: Position Encoding in PEVLM
As described in observations, similar to the case in LLMs, applying parallel-encoding-based methods in VLMs suffers from the misalignment of attention weight distributions when compared to Full-Attention. However, introducing additional hyperparameters to adjust attention weights for better alignment is not ideal for efficient deployment. Therefore, inspired by Block Attention (Ma, Wang, and Tian 2025) and TurboRAG (Lu et al. 2024), we retain sequential position embeddings as illustrated in Figure 3.
Due to retaining the sequential position embeddings, PEVLM does not extend the model’s inherent capacity to handle longer video contexts, unlike some other parallel encoding-based approaches. For instance, LongVILA-7B-256f is generally trained to process up to 256 frames, and feeding longer videos may lead to accuracy degradation. Although PEVLM accelerates inference, it still processes 256 frames, and longer inputs may still result in degraded performance. However, this is acceptable, as our primary objective is to improve prefilling efficiency.
@@
The main objective of this experiment is to analyze the impact of PEVLM on accuracy in long-video understanding tasks. We have chosen the LongVideoBench (Zhang and Wang 2024), VideoMME (Fu et al. 2024), EgoSchema (Mangalam, Akshulakov, and Malik 2023), and MVBench (Li et al. 2024a) as benchmarks, which are designed to comprehensively evaluate multimodal models in video understanding. We have carried out experiments using LLaVA-Video (Zhang et al. 2024), LongVILA (Chen et al. 2024), and Qwen2.5-VL (Bai et al. 2025) as test models. To minimize any adverse effects on accuracy, all evaluations were performed using bf16 precision.
As shown in Table 1, we compared the accuracy of Full-Attention, Block-Attention, APE, Star-Attention, and PEVLM in the datasets. Since the other methods are finetune-free, we did not finetune the weights in the Block Attention test. For APE, to accommodate the needs of large-scale deployment, we tested with T=0. Regarding the sink block (shared prefix), we selected the system prompt by referring to the code provided by APE. For Star-Attention, we implemented its equivalent algorithm on a single node, with the anchor size set equal to the block size, as the configuration in the Block-Attention paper (Ma, Wang, and Tian 2025). The block size was uniformly set to 4096 across all algorithms. Smaller block sizes resulted in reduced accuracy for all methods, while larger blocks significantly degraded inference performance. For PEVLM, we set the sink size and block size to about 16 frames, which corresponds to approximately 4k tokens for Qwen2.5-VL and LongVILA, and about 3k tokens for LLaVA-Video.

### v3

@@ Introduction
In recent years, Vision-Language Models (VLMs) have become a central research focus at the intersection of computer vision and natural language processing. These models have demonstrated impressive performance across a wide range of multimodal understanding and generation tasks (Alayrac et al. 2022; Li et al. 2020; Radford et al. 2021; Zhu et al. 2023). As their capabilities continue to improve, VLMs are being applied in increasingly complex domains, including robotics (Black, Lee et al. 2024; Cheang, Huang et al. 2024; Prasad, Singh et al. 2024), autonomous driving (Gao, Lin et al. 2024; Hu, Wu et al. 2023; Wang, Zhao et al. 2024), and healthcare (Liu, Zhang et al. 2024). These application scenarios often demand processing longer video inputs, placing significant strain on computational resources.
A key obstacle to applying VLMs to long-video inputs is the quadratic complexity of transformer attention during the prefilling stage (Vaswani et al. 2017; Beltagy, Peters, and Cohan 2020). This computational burden limits scalability in tasks such as perception, temporal reasoning, and action planning, which are fundamental to domains like robotics, autonomous driving, and healthcare. As a result, optimizing the prefilling phase is critical to enabling efficient long-video processing with VLMs.
To address this issue, a widely adopted technique in large language models (LLMs) is the parallel encoding mechanism (Li et al. 2024b; Liu et al. 2024; Ma, Wang, and Tian 2025; Ratner et al. 2023; Yang, Chen, and Chen 2025; Yen, Gao, and Chen 2024; Lu et al. 2024). In this approach, the input context is divided into multiple blocks, with each block independently encoded into Key-Value (KV) states. This design prevents tokens in different blocks from attending to one another, thereby reducing the computational complexity from $O(L^{2})$ to $O(L)$. Moreover, parallel encoding alleviates the issue of “lost in the middle” by reducing the number of tokens participating in the softmax operation and can even achieve accuracy surpassing that of Full-attention in long-context scenarios (Liu et al. 2023; Yang, Chen, and Chen 2025; Veličković et al. 2025).
@@
In addition, there are several widely discussed methods such as Native Sparse Attention (Yuan et al. 2025) from Deepseek and MoBA (Lu et al. 2025), which adopt similar algorithmic principles and achieve acceleration effects comparable to parallel encoding. However, both approaches involve modifications to the model architecture and thus require retraining from scratch. This poses a challenge for many inference platforms, which typically do not have control over users’ model preferences. The issue is particularly pronounced in domains such as autonomous driving, where a reliable models are often continuously trained and iterated over years. Consequently, we focus on finetune-free performance optimization methods to ensure compatibility and ease of deployment.
Given all of these, a key question naturally arises: Can existing methods be directly applied to VLMs to accelerate inference and deployment pipelines? In Table 1, we evaluate several models with existing parallel encoding methods on LongVideoBench (Zhang and Wang 2024), VideoMME (Fu et al. 2024), EgoSchema (Mangalam, Akshulakov, and Malik 2023) and MVBench (Li et al. 2024a) benchmarks. Although existing parallel encoding methods demonstrate certain effectiveness on shorter videos, we observe varying degrees of accuracy degradation across different models when applied to longer videos. This performance drop is particularly pronounced in models like Qwen2.5-VL, which are more sensitive to spatial and temporal information in videos—sometimes resulting in completely empty or garbled outputs in long-video scenarios, with accuracy dropping to 0.0%. To investigate this issue, we analyze the attention weight distributions under parallel encoding and identify a misalignment between Full-Attention and parallel encoding. Existing methods developed for LLMs fail to effectively address this misalignment in VLMs. Building upon this observation, we conducted a series of studies and experiments and ultimately propose a novel attention mechanism for VLMs, termed Parallel Encoding for Vision-Language Models (PEVLM). PEVLM is a fine-tuning-free method specifically designed to accelerate the prefilling stage in long-video processing with VLMs. Our contributions involve:
@@
  • We propose PEVLM to recover the accuracy of parallel encoding by applying three targeted alignment steps: (i) segmenting contexts into blocks by video frames rather than tokens; (ii) using the system promots and the initial frames of the video as Sink Block for all blocks to avoid the duplication of abnormal distribution of initial tokens; (iii) preserving the sequential position embeddings instead of resuing position embeddings across blocks. With these alignment strategies, PEVLM achieves higher accuracy than existing parallel encoding methods, and reaches 99.57% to 104.80% of the accuracy of full attention on different models.
  • PEVLM reduces the computational complexity of the attention mechanism from $O((T\times N)^{2})$ to $O(T\times N)$, which leading to a significant improvement in prefilling efficiency. (i) For 100k (text&video) token contexts prefilling, the attention layer achieves a 7.47× speedup, while the end-to-end 2.58x speedup without compromising generation quality. This highlights the practical viability and superiority of PEVLM in cloud deployment scenarios. (ii) Under fixed latency constraints, PEVLM increases accuracy from 23.26% to 61.03%, showcasing its critical value for latency sensitive applications.
@@
Figure 1: We collected the distribution of attention weights under different scenarios and computed their moving averages for better observation. The top row shows the attention weights for all tokens, while the bottom row shows those for the visual tokens. (a) Full-Attention: The textual tokens and the initial video frames receive the highest attention weights, whereas the remaining video frames are assigned relatively lower attention weights. (b) Parallel Encoding (Reusing Position Embedding): The overall attention weights distribution resembles Full-Attention, with higher attention weights for the textual tokens and initial video frames. While compared to Full-Attention, the initial video frames receive even higher attention weights, and the remaining video frames exhibit more uniform and lower attention weights. (c) Parallel Encoding (Sequential Position Embedding): The overall attention weights distribution remains similar to (a) and (b). However, the attention weight distribution within the video frames more closely aligns with (a)).
@@
(i) As shown in Figure 1(a), some initial text and visual tokens receive high attention scores, consistent with the observations reported in (Lab and AI 2023; Sun et al. 2024; Kang et al. 2025). This suggests that it may be necessary to include early video frames when selecting sink tokens.
@@
These observations yield two key insights for applying parallel encoding to VLMs: (i) the shared sink block in parallel encoding should include early video frames in addition to textual prompts, as they function as important attention sinks; (ii) maintaining the sequential position embeddings of visual tokens is essential for preserving temporal consistency, and reusing position embeddings across blocks could disrupt temporal alignment. These findings underscore the need for a VLM-specific parallel encoding strategy, which we introduce in the next section.
@@
With the observation in the last section, we design PEVLM to improve computational efficiency. PEVLM enables a seamless shift from Full-Attention to parallel encoding without requiring training while maintaining most of the model’s capabilities. Our approach adaptively aligns the distribution of attention weights between Full-Attention and parallel encoding, thereby boosting efficiency and performance.
@@
Before introducing PEVLM, we briefly review key aspects of the standard attention mechanism as a foundation. The computation formula for standard softmax attention is:
@@
Attention Sink. The attention sink phenomenon was first identified by StreamingLLM (Lab and AI 2023), which observed that one or more tokens at the beginning of the input consistently receive higher attention scores. APE (Yang, Chen, and Chen 2025) and Star-Attention (Liu et al. 2024) have demonstrated that sharing a common attention sink across parallel encoding blocks significantly enhances the accuracy of the model. Some studies also show that the attention sink phenomenon exists in VLMs but is more distributed among visual tokens (Kang et al. 2025).
@@
Position Embedding. Position embedding provides essential information that distinguishes the position and order amongst tokens in a sequence. For VLMs, contextual positional information is more critical. In particular, for some VLM models like Qwen2.5-VL (Bai et al. 2025), their position embeddings also incorporate spatio-temporal information between video frames, making position embeddings increasingly essential for VLMs.
Computational Complexity. In standard softmax attention mechanisms, the total number of operations in the prefilling phase is given by
@@
Position Encoding
Figure 3: Position Encoding in PEVLM
As described in the Observations section, similar to the case in LLMs, applying parallel encoding methods in VLMs suffers from the misalignment of attention weight distributions when compared to Full-Attention. While prior work has attempted to address this issue by introducing additional hyperparameters to realign attention scores (Yang, Chen, and Chen 2025), such approaches are not ideal for efficient deployment. To address this, PEVLM draws inspiration from Block Attention (Ma, Wang, and Tian 2025) and TurboRAG (Lu et al. 2024), and retains sequential position embeddings, as shown in Figure 3. In contrast to these methods, which dynamically adjust positional embeddings, often to accommodate context reuse through sequence reordering, PEVLM forgoes such updates. Since video frames are processed in their original temporal order, PEVLM directly applies the sequential position embedding to each visual token within context blocks, without modifying their positions after query concatenation.
By preserving sequential position embeddings, PEVLM maintains attention weight distributions that closely resemble those of Full Attention. However, unlike some parallel encoding methods that aim to extend context capacity, this design choice also constrains PEVLM’s ability to scale to longer video sequences. For instance, LongVILA-7B-256f is typically trained on sequences up to 256 frames; attempting to process longer inputs may lead to performance degradation. Although PEVLM achieves significant acceleration of inference, it is similarly limited to 256-frame inputs. Nonetheless, this trade-off is acceptable as our primary goal is to enhance inference efficiency rather than to expand context length.
Model Method MVBench EgoSchema VideoMME LongVideoBench (Val) Avg.
[…]
Qwen2.5-VL- Full Attn 68.47% 57.00% 74.56% 56.72% 72.85% 55.33% 64.16%
7B-Instruct Block Attn 66.03% 47.40% 65.22% 50.39% 70.64% 47.85% 57.92%
Frames: 768 APE(T=1.0) 66.03% 57.20% 50.56% 1.06% 54.02% 3.38% 38.71%
Tokens: 100k Star Attn 60.19% 19.80% 3.00% 0.00% 49.03% 0.00% 22.00%
PEVLM 68.44% 61.00% 74.33% 58.83% 74.52% 57.58% 65.78%
LongVILA- Full Attn 61.92% 57.00% 66.33% 49.39% 65.10% 47.44% 57.86%
7B-256f Block Attn 57.81% 59.00% 64.89% 51.44% 61.77% 48.05% 57.16%
Frames: 256 APE(T=1.0) 60.00% 60.60% 68.56% 51.61% 63.71% 46.93% 58.57%
Tokens: 66k Star Attn 63.39% 61.60% 70.11% 54.17% 63.43% 47.44% 60.02%
PEVLM 63.53% 61.40% 69.44% 54.33% 65.93% 49.18% 60.64%
LLaVA-Video- Full Attn 60.67% 58.20% 75.33% 59.99% 72.02% 55.53% 63.62%
7B-Qwen2 Block Attn 56.86% 51.60% 67.56% 55.50% 63.99% 52.36% 57.98%
Frames: 128 APE(T=1.0) 58.00% 57.60% 72.22% 58.19% 67.87% 54.61% 61.42%
Tokens: 26k Star Attn 59.31% 57.80% 72.89% 58.71% 68.98% 54.92% 62.10%
PEVLM 60.00% 58.20% 74.33% 59.99% 72.02% 55.53% 63.35%
@@
The computational formulation of PEVLM is defined as:
$\text{Attn}s=f(Q_{s},K_{s},V_{s}),$ (3)
@@
where $s$ denotes the Sink Block, $q$ denotes the Question Block, and $c_{i}$ denotes the $i$-th Context Block. Symbols $Attn_{s}$, $Attn_{c_{i}}$, and $Attn_{q}$ represent the attention outputs for the corresponding blocks. The function $f(\cdot)$ corresponds to the standard softmax attention defined in Equation (1), with $Q$, $K$ and $V$ denoting the query, key, and value matrices respectively.
The total computational operation (OP) count of PEVLM is:
$\text{OP}_{\text{PEVLM}}=\text{OP}_{\text{Sink}}+N\times\text{OP}_{\text{ContextBlock}}+\text{OP}_{\text{Quest}},$ (6)
$\text{OP}_{\text{Sink}}=2HS^{2},$ (7)
$\text{OP}_{\text{ContextBlock}}=2HB(S+B),$ (8)
$\text{OP}_{\text{Quest}}=2HQL,$ (9)
where $S$ denotes the sink block size, $N$ denotes the context block number, $B$ denotes the context block size, $Q$ denotes the query block size and $H$ is the hidden size. $L$ is the total token number and $L=S+N\times B+Q$.
@@
As the context block size is a static number, the computational complexity of PEVLM is $O(L)$.
Block Size
As analyzed in the previous section, the number of frames selected for the Sink Block depends primarily on the distribution of attention weights. The optimal strategy is to include all tokens with significantly higher attention scores at the beginning of the video in the Sink Block. However, on the one hand, the number of frames containing visual tokens with a high attention score varies across different videos; on the other hand, as shown in Equation 10, a larger sink block size leads to higher latency, resulting in a trade-off between accuracy and efficiency. A similar trade-off also exists in determining the context block size. We will further analyze this in the following sections.
@@
This experiment aims to evaluate the impact of PEVLM on accuracy in long-video understanding tasks.
@@
We adopt LongVideoBench (Zhang and Wang 2024), VideoMME (Fu et al. 2024), EgoSchema (Mangalam, Akshulakov, and Malik 2023), and MVBench (Li et al. 2024a) as benchmarks, which are designed to comprehensively evaluate multimodal models in video understanding. All evaluations are conducted using the lmms-eval toolkit (Li* et al. 2024). We test on LLaVA-Video (Zhang et al. 2024), LongVILA (Chen et al. 2024), and Qwen2.5-VL (Bai et al. 2025) as representative models. To minimize accuracy loss, all experiments are conducted with bf16 precision.
We compared the accuracy of Full-Attention, Block-Attention, APE, Star-Attention, and PEVLM in the datasets. Since the other methods are finetune-free, we did not finetune the weights in the Block Attention test. For APE, we conducted experiments with a temperature setting of $T=1.0$. This choice was motivated by two considerations. First, lowering the temperature further leads to additional accuracy degradation. As previously observed, parallel encoding tends to produce lower attention weight distributions over context blocks in VLMs, rather than higher. Therefore, decreasing the temperature further skews the attention distribution away from that of Full-Attention, exacerbating the accuracy loss. Second, using $T=1.0$ is favorable for large-scale deployment. Regarding the sink block (shared prefix) size, we selected the system prompt by referring to the code provided by APE. For Star-Attention, we implemented its equivalent algorithm on a single node, with the anchor size set equal to the context block size, following the setup in the Block-Attention paper (Ma, Wang, and Tian 2025). All methods use a context block size of 4096 tokens. Smaller context block sizes resulted in reduced accuracy for all methods, while larger blocks significantly degraded inference performance. For PEVLM, we set the sink and context block size to 16 frames, which corresponds to approximately 4k tokens for Qwen2.5-VL and LongVILA, and about 3k tokens for LLaVA-Video.